OpenStudy (anonymous):

Suppose a bank loans $4,500 at an interest rate of 5.5% compounded annually. Find the total amount of interest owed on the loan after 6 years. A. $247.50 B. $1485.00 C. $1704.79 D. $2475.00

OpenStudy (anonymous):

I think A. if not right sorry lol

OpenStudy (cj49):

Remember the formula for compound interest is A=P(1+r/100)^t Where A stands for Amount, P stands for Principal/Loan, R stands for Rate and T stands for time

OpenStudy (anonymous):

Im thinking its C for some reason

OpenStudy (anonymous):

*calculating*

OpenStudy (cj49):

4500(1+5.5/100)^6=?

OpenStudy (cj49):

6204-4500=?

OpenStudy (anonymous):

I knew it! it's C!

OpenStudy (cj49):

lol....yes thats the ans